OpenStudy (anonymous):

ok solve using 7x+4y=-13 -9x+y =29

OpenStudy (anonymous):

x = 3, y = -8.5

OpenStudy (anonymous):

wrong y

OpenStudy (gw2011):

Here is one way of solving it: Multiply the second equation by -4 to eliminate y and then you get: 7x+4y=-13 36x-4y=-116 Add the two above equations and you get: 43x=-129 x=-3 Substitute x=-3 into one of the equations to solve for y: 7x+4y=-13 7(-3)+4y=-13 -21+4y=-13 4y=8 y=2 Therefore, x=-3 and y=2

OpenStudy (anonymous):

ty solved